OpenStudy (anonymous):

x+y-5z=-5 y-2z=14 4y-2z=8

OpenStudy (anonymous):

find the X,Y,Z

OpenStudy (anonymous):

do u want to solve it by elimination?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply second equation by -1 then add it to third equation

OpenStudy (anonymous):

i dont think i did it correct @mukushla

OpenStudy (anonymous):

@saifoo.khan can you help ?

OpenStudy (saifoo.khan):

Which method you know good? Elimination or substitution? @sharee2012

OpenStudy (anonymous):

substitution i think

OpenStudy (saifoo.khan):

x+y-5z=-5 y-2z=14 4y-2z=8 taking equation 2. solving for y. y -2z = 14 y = 14 + 2z Now use this value and insert it in equation 3 let's say. then you will get the value of z. 4y-2z=8 y(14+2z) -2z = 8 Solve..

OpenStudy (anonymous):

14?

OpenStudy (saifoo.khan):

Typo in the last line: 4*(14+2z) -2z = 8

OpenStudy (anonymous):

y(14+2*3)-2*3=14

OpenStudy (saifoo.khan):

Solve this: 4*(14+2z) -2z = 8

OpenStudy (anonymous):

74?

OpenStudy (saifoo.khan):

im getting z = -8